Mir Status Report 20 September 1997

Date: Saturday, September 20, 1997// MD 46/128

MIR STATION SYSTEM/SUBSYSTEM ACTIVITIES

Vacuum valve #2 (BVK-2) on Vozdukh failed at DMT 15:20. It had been operating for a total of 30,000 hours, well over its service life of 20,000 hours. The crew was instructed tonight to begin working procedures to reactivate Vozdukh on the remaining two valves. The SFD informed us that the valve would be replaced tomorrow. The crew also performed troubleshooting procedures on the urine water recovery system SRV-U, which has been unable since yesterday to drain the urine residue from the distillation process into the residue EDV. These procedures included attempts to drain the residue through various connectors into another EDV, but were unsuccessful in determining the location of the clog. The SFD informed us that they may have the crew replace the rotary pump and other equipment to fix the problem.

MIR PAYLOAD ACTIVITIES

Greenhouse Daily Ops (FE-1); CPA; Beetle Power Status Check
PUP check-out; MIPS check-out; Private Family Conference
OTHER ACTIVITIES: Physiolab

Experiment Hardware / PUP and MIPS checkout - Mike informed the ground crew that Priroda had not been powered up yet , therefore he would have to slip the PUP and MIPS checkout to a later date. Further discussions with the shift flight director revealed that the reason the Priroda had not been powered was due to the ground's inability to issue commands to the Priroda's Master Computer. It is believed that the reason for this is condensation in connectors. There is a concern that issuing commands to route power to the PUPs may result in an erroneous command being issued by the Master Computer due to condensation. The computer must be deactivated before the connectors can be cleaned and the specialists needed to issue the commands for deactivation have gone home for the weekend. We have emphasized the absolute requirement that the Priroda be fully functional and the PUP, MIM and BTS systems be fully checked out prior to launch of STS-86. The shift flight director has made every effort to get in contact with the Priroda systems specialists, but so far has had no success. There is one gentleman who has been located who is a general module expert who is tried to give the crew commands to shut down a portion of the Master Computer and that was unsuccessful. Further commands will be issued by the ground this evening.

MiSDE Sensors - Mike reported that he had checked all the MiSDE sensors. All the uniaxial sensors were oriented incorrectly and had to be reoriented. Mike stated that it took between 45 minutes and one hour for each sensor. He will send down a written summary.

Stowage/Transfer -- Mike had questions on the work messages sent up yesterday (u0919a_w.txt) concerning the packing of the cushion packing assemblies. He would like to pack those earlier than the handover. He feels that he has more than enough CTB for what needs to be packed and wants to pack the foam accessories in CTBs. He also questioned the return of glovebox cushions assembly. Mike feels that this is useful and wanted to know why this is being returned.

Another item in the list which he questioned was the Mir Safety Tether. He was unclear how many this was referring to and whether they were being exchanged. Information was obtained from NASA EVA personnel and will be forwarded to Mike.

SMP/Medical - Nothing to report

Environmental Parameters (Previous 24 hour period):
Parameter Min. Max. Nominal Range
Pressure (mmHg) BB 617 618 660-850
ppO2 (mmHg) 151 152 140-200
ppCO2 (mmHg) 4.4 5.1 up to 8.0
ppH20 (mmHg) 13.6 14.8 8-14
Humidity (percentage) 56 60 30-75
Temperature (C) (BB) 25.2 25.8 18-28

PLANNING/MIR 24/NASA 5 MD 47/129
Daily Exercise (Treadmill); A/G Comm with France
NASA MIR PROGRAM / Mir 24 / NASA 5 Status
Date: Sunday, September 21, 1997 // MD 47/129

MIR STATION SYSTEM/SUBSYSTEM ACTIVITIES

Vozdukh was activated using two valves today. It can keep up with most of the CO2 production but must be supplemented by LiOH cans during peak periods (such as exercise). Maintenance activities should be later in the week.

Maintenance on the SRV-U should be performed tomorrow. All other systems are in the same state as yesterday.

OTHER ACTIVITIES
Priroda activation; PFCs for CDR and FE-1; TM Connector Check Mike was asked to change BC values to 40. He reported that there are approximately 9-10 space seeds sprouting and 15 earth seeds. Tomorrow he will provide a row by row count.

Experiment Hardware / PUP and MIPS checkout - PUP A, PUP B, and MIPS were successfully checked with one minor problem. The LED on J9 on PUP A did not light up. Whether this signifies a broken LED or no power to that port is not known at this time. Further analysis of that one port will be needed to determine the problem.

Stowage/Transfer - MOST gave Mike a description of the CCD that is in the prepack list. Mike was also given a go to pack any remaining CGEL items except the three pieces of the photo setup that are needed for NASA 6.

SMP/Medical: CPA data (taken in Priroda on 9/20):
HCl = 0.1; HCN = 0.1; CO = 18
Mike was asked to repeat the CPA readings today in the Priroda module due to the CO reading.

Mike participated in the first day of the physical training cycle with recording on the `ZY'.

In preparation for the arrival of STS-86, the entire Mir crew will sleep shift tonight. Sleep starts at 23:00 DMT tonight and wake up is at 14:00 DMT on Monday, 22 September.

Environmental Parameters (Previous 24 hour period):
Parameter Min. Max. Nominal Range
Pressure (mmHg) BB 617 620 660-850
ppO2 (mmHg) 151 153 140-200
ppCO2 (mmHg) 4.3 6.7 up to 8.0
ppH20 (mmHg) 13.5 14.9 8-14
Humidity (percentage) 53 58 30-75
Temperature (C) (BB) 24.9 26.5 18-28

COMMUNICATIONS SUMMARY
A/G with Consultants Group
1012-1013 Greenhouse
1145-1149 Greenhouse, packing, pre-launch requirements
1457-1504 Comments for next mission
1919-1922 PUPS, MIPS
2047-2048 Personal

Packet -- The form 24, a weekly overview, and one personal file was submitted for uplink. One work file and two personals were received from Mike.

PLANNING
MIR 24/NASA 5 MD 48/130
BTS Checkout
MIM Power Up
Daily Exercise (Treadmill) with recording on `ZY'
PFC

  • submit to reddit


Loading



MILITARY SPACE USA, 11TH JUNE TO 12TH JUNE 2019, LOS ANGELES, USA
49th International Conference on Environmental Systems - ICES 2019




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on September 20, 1997 8:43 PM.

Mir Status Report 19 September 1997 was the previous entry in this blog.

Mir Status Report 20 September 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.